# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v7

@@ 2.5 Noisy Information
In fact, for both information $I$ and its noisy information $\tilde{I}$, it is easy to verify its recoverability or recoverable reducibility from the perspective of enabling mapping. However, in real scenarios, it is often difficult to obtain noise-free information $I$, and one can only study using a noisy version $\tilde{I}$. Moreover, the recoverable state of $\tilde{I}$ is not the same as that of $I$, which is the fundamental reason for the skepticism about the recoverability of information. But under the condition where the noiseless information $I$cannot be directly obtained, we can approximate the ideal state of $I$ using $\tilde{I}$ along with the relationships between $S_{o}(o,T_{h}),S^{-}_{o}(o,T_{h})$, and $S_{o}^{+}(o,T_{h})$. This is the practical significance of introducing the concept of noisy information.
Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenario[24]. In this case, Shannonassumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
@@
where $\widetilde{S_{I_{c}}}(I_{s},T_{h})=S_{I_{s}}(I_{s},T_{h})\backslash S_{I_{s}}^{-}(I_{s},T_{h})\cup S_{I_{s}}^{+}(I_{s},T_{h})$.
In this way, we use the definition of noisy information to explain Shannon’s model of noisy channels in[24], demonstrating that OIT is entirely applicable to the complex scenarios of information research under noisy conditions.
@@ 3 Meta-Framework for Machine Learning Theory (MLT-MF)
The core of machine learning is the model, which is a typical finite automaton and also an information system. Therefore, every learning process and the subsequent testing or application process rely on the information within it, which can generally be summarized and divided into four main steps of information realization and operation.
Step 1: The implementation of training information $It$, in which the ontological state of $It$ is enabled and mapped onto the model interface as its carrier;
Step 2: The original information $Io$ within the model—endowed with learning capability—is fused with the training information $It$, and optimized to yield the updated information $Iu$, which possesses both the ability to answer specific tasks and the capacity for further learning;
Step 3: The implementation of user input information $Iq$, where the ontological state of $Iq$ is likewise enabled and mapped onto the model interface as its carrier;
Step 4: The updated information $Iu$ within the model processes the input $Iq$, outputs the response information $Ir$ according to user requirements, and simultaneously undergoes further optimization through learning, resulting in $Iv$.
@@ 3.1 Formal Modeling and Realization of Training Information
According to the general formal definition of information, the formal representation of training information implementation is straightforward:
$It=(ot,Tt_{h},S_{ot},ct,Tt_{m},S_{ct})$ (3.1)
$S_{ot}(ot,Tt_{h})=\{\varphi t|\varphi t\text{ is a well-formed formula concerning }ot\text{ and }Tt_{h}\}$ (3.2)
$S_{ct}(ct,Tt_{m})=\{\psi t|\psi t\text{ is a well-formed formula concerning }ct\text{ and }Tt_{m}\}$ (3.3)
[…]
Therefore, it is necessary to introduce a model training data interface $ct$—akin to hardware interfaces or software APIs—as the carrier for $It$, so that $S_{ot}(ot,Tt_{h})$ can be ingested in an engineering-compatible manner. The reflection time set $Tt_{m}$ denotes the period during which the carrier reflects $It$, while the carrier reflection set $S_{ct}(ct,Tt_{m})$ comprises specific, complete, and operational realizations of $It$ on $ct$ at $Tt_{m}$. Each element $\psi t$ is a well-formed formula enabled by, and logically/semantically corresponding to, some $\varphi t$.
Continuing the example above, if $S_{ot}(ot,Tt_{h})$ is enabled and mapped to $ct$ at time $T_{3}$, we might express
$S_{ct}(ct,T_{3})=\{\psi t\}=\{(\text{TrainTable}[\text{user}_{id}=X][\text{time}=T_{1}]=\text{Active})\land$
[…]
where “Active” and “Tgt60” are defined in the training table as “user is active” and “device temperature exceeds 60°C,” respectively. Although the form of $\psi t$ may appear simple, it fully reflects the semantic content of $\varphi t$. In practice, $\psi t$ may take the form of tables, vectors, JSON objects, API parameters, or other formal representations of engineering data interfaces or hardware device states, and may include additional data constraints or format checks to ensure the validity and robustness of $S_{ct}(ct,T_{3})$.
It is crucial to emphasize that $S_{ot}(ot,Tt_{h})$ is only realized as $S_{ct}(ct,T_{3})$ via the enabling mapping of $It$. The fact that $S_{ot}(ot,Tt_{h})$ satisfies Postulate 5 also implies that $S_{ct}(ct,T_{3})$ satisfies Postulate 5. This mapping embodies both the structural-semantic correspondence from the former to the latter and the causal relationship—where the ontological states are the cause and the carrier states are the effect, requiring expenditure of human, temporal, and energy resources. Such an enabling mapping allows the knowledge, facts, and rules formalized within the ontology to be incorporated into the model in a structured, standardized, and machine-readable fashion, thereby becoming usable for subsequent model learning.
In practical engineering, the realization of training information often needs to consider noisy information $\widetilde{It}$ due to discretization, attribute pruning, and noise accumulation. However, for theoretical analysis, we currently focus only on the ideal case of no noise, leaving the study of complex noisy conditions for further research.
@@ 3.2 Formal Modeling and Implementation of Model Learning
From the perspective of information flow and transformation, the learning process of an artificial intelligence model M can be viewed as the optimization of its existing information $Io$ through learning from the training information $It$, resulting in the updated information $Iu$. The preceding section has already detailed the modeling and realization of training information $It$. We now turn our attention to the modeling and implementation of $Io$ and $Iu$ within M. By analogy with equations (3.1-3.3) for $It$, we can define $Io$ and $Iu$ by substituting the respective subscript $t$ with $o$ and $u$. It is important to note that, since $Io$ and $Iu$ represent the information present in M before and after learning, their carriers $co$ and $cu$ are effectively identical—in other words, both can directly be denoted by M, or more precisely, by the computational system that executes M:
$co=cu=M$ (3.4)
Within the learning process of M, the most crucial component is the existing information $Io$, which, in addition to encapsulating knowledge, propositions, facts, or samples (as with $It$), is distinguished by its possession of learning rules—i.e., the rules by which it is optimized into updated information through learning.
Definition 8 (State with Learning Capability): Let the states $Sx(x,T_{x})$, $Sy(y,T_{y})$, and $Sz(z,T_{z})$ be interpretations $E_{x}$, $E_{y}$, and $E_{z}$, respectively, of three sets of well-formed formulas within the same formal system $\mathcal{L}$, corresponding to object $x$ over time interval $T_{x}$, object $y$ over $T_{y}$, and object $z$ over $T_{z}$, all satisfying Postulate 5. We introduce the following specific functions and predicates in $\mathcal{L}$:
  • Function: $f_{learn}^{2}(X,Y)$ denotes the well-formed formula resulting from the learning of set $X$ from set $Y$. Moreover, when both $X$ and $Y$ satisfy the logical consistency requirement of Postulate 5, the function $f_{learn}^{2}(X,Y)$ also satisfies this logical consistency requirement.
  • Predicate: $\text{IsSubsetOf}^{2}(X,Y)$ indicates that set $X$ is a subset of set $Y$; $\text{Learnable}^{2}(X,Y)$ indicates that state set $X$ can learn from state set $Y$; $\text{IsFormulaOf}^{2}(x,X)$ indicates that $x$ is a well-formed formula within state set $X$; and $\text{Eq}^{2}(x,y)$ denotes that $x$ is equal to $y$.
Then suppose $\varphi x\in Sx(x,T_{x})$, $\sup(T_{x},T_{y})\leq\inf(T_{z})$, and
$\varphi x=$ $(\forall\Phi x)(\forall\Phi y)(\text{IsSubsetOf}^{2}(\Phi x,Sx(x,T_{x}))\land\text{IsSubsetOf}^{2}(\Phi y,Sy(y,T_{y}))\land$
$\text{Learnable}^{2}(\Phi x,\Phi y))\rightarrow\exists\varphi z(\text{IsFormulaOf}^{2}(\varphi z,Sz(z,T_{z}))\land\text{Eq}^{2}(\varphi z,f_{learn}^{2}(\Phi x,\Phi y)))$ (3.5)
we say that $Sx(x,T_{x})$ has the capability to learn from $Sy(y,T_{y})$ and optimize into $Sz(z,T_{z})$; equivalently, we say that the interpretation $E_{x}$ has the capacity to learn from $E_{y}$ and optimize toward $E_{z}$.
The formula $\varphi x$ essentially constitutes the learning rule for the state $Sx(x,T_{x})$, indicating that as long as $Sy(y,T_{y})$ meets the learning conditions of $Sx(x,T_{x})$, it is possible to augment $Sz(z,T_{z})$ with new well-formed formulas $\{\varphi u|\varphi u=f_{learn}^{2}(\Phi x,\Phi y),\Phi x\subseteq Sx(x,T_{x}),\Phi y\subseteq Sy(y,T_{y})\}$, thus rendering $Sz(z,T_{z})$ an optimized extension of $Sx(x,T_{x})$. Furthermore, given the logical consistency of $f_{learn}^{2}(X,Y)$, if $Sx(x,T_{x})$ and $Sy(y,T_{y})$ satisfy Postulate 5, then $Sy(y,T_{y})$ also satisfies Postulate 5.
On the other hand, the inheritance of learning capability is also a fundamental property of model learning. In fact, to inherit the learning capability of $Sx(x,T_{x})$, it suffices to retain the learning rule $\varphi x$ within $Sz(z,T_{z})$, with only partial adaptive substitution of its symbols. Specifically, for $\varphi z\in Sz(z,T_{z})$:
$\varphi z=($ $\forall\Phi z)(\forall\Phi y)(\text{IsSubsetOf}^{2}(\Phi z,Sz(z,T_{z}))\land\text{IsSubsetOf}^{2}(\Phi y,Sy(y,T_{y}))\land$
$\text{Learnable}^{2}(\Phi z,\Phi y))\rightarrow\exists\varphi w(\text{IsFormulaOf}^{2}(\varphi w,Sw(w,T_{w}))\land$
$\text{Eq}^{2}(\varphi w,f_{learn}^{2}(\Phi u,\Phi y)))$ (3.6)
This ensures that $Sz(z,T_{z})$ can learn from $Sy(y,T_{y})$ and optimize into $Sw(w,T_{w})$. Thus, we establish:
Proposition 3 (Inheritance of Learning Capability): If the state set $Sx(x,T_{x})$ is capable of learning from $Sy(y,T_{y})$ and optimizing into $Sz(z,T_{z})$, then $Sz(z,T_{z})$ always inherits the learning capability of $Sx(x,T_{x})$.
Definition 9 (Realization of Model Learning): For an AI model M with original information $Io$, training information $It$, and updated information $Iu$, if all three are recoverable, and the ontological state $S_{oo}(oo,To_{h})$ can learn from $S_{ot}(ot,Tt_{h})$ and be optimized into $S_{ou}(ou,Tu_{h})$, and the carrier state $S_{co}(co,To_{m})$ can likewise learn from $S_{ct}(ct,Tt_{m})$ and be optimized into $S_{cu}(cu,Tu_{m})$, then M is said to possess the capability to realize the learning and optimization of $Io$ towards $It$, resulting in $Iu$—in short, M has information learning capability.
According to (3.4), we may also write $S_{co}(co,To_{m})=S_{M}(M,To_{m})$, $S_{cu}(cu,Tu_{m})=S_{M}(M,Tu_{m})$, indicating that the learning process is essentially an optimization of the state of model M.
In summary, we have theoretically internalized the model’s learning capability as a property of its states. These rules are independent of the algorithmic structure or scale of the model, thereby providing a unified and general formal framework for the analysis of various types of models.
As an example, let $Mn$ be a neural network model. According to common knowledge, the information in $Mn$ is $Io=\langle oo,To_{h},S_{oo},Mn,To_{m},S_{co}\rangle$, where $oo$ represents the model in the designer’s mind, $To_{h}$ is the time set during the design of $Mn$.
$S_{oo}(oo,To_{h})=\Phi o=\{Params(\theta)\land Architecture(L_{1},\ldots,L_{n})\land Activation(A)\}$ (3.7)
Here, $Params(\theta)$ denotes the predicate representation of the parameter set $,\theta=\{W_{1},b_{1},\ldots$$,W_{n},b_{n}\}$; $Architecture(L_{1},\ldots,L_{n})$ denotes the predicate representation of the network layer structure; and $Activation(A)$ denotes the predicate representation of the activation function set $A=\{a_{1},\ldots,a_{n}\}$.
The carrier for $Io$ is the neural network model $Mn$ itself, with $To_{m}$ representing the period after its development, and $S_{co}(co,To_{m})$ being the physical instantiation of (3.7) within $Mn$.
The training information for $Mn$ is $It=\langle ot,Tt_{h},S_{ot},ct,Tt_{m},S_{ct}\rangle$, where $ot$ is a set of real-world objects, $Tt_{h}$ is the time set during which the state $S_{ot}(ot,Tt_{h})$ occurs, and
$S_{ot}(ot,Tt_{h})=\Phi t=\{Batch(X,y)\land DimMatch(X,L_{1})\land Differentiable(L)\}$ (3.8)
where $Batch(X,y)$ expresses the predicate for input data $X\in\mathbb{R}^{b\times k}$ and labels $y\in\mathbb{R}^{b\times k}$; $DimMatch(X,L_{1})$ expresses that the input data $X$ matches the dimensions of the input layer $L_{1}$; and $Differentiable(L)$ expresses the differentiability of the loss function $L$ with respect to parameters $\theta$.
The definition of the predicate $Learnable^{2}$ is as follows:
$Learnable^{2}(\Phi o,\Phi t)=$ $Dimensionality(\Phi o,\Phi t)\land ActivationFeasible(\Phi o,\Phi t)\land$ (3.9)
$GradientExists(\Phi t)$
where $Dimensionality(\Phi o,\Phi t)=DIM(X,b\times d)\land DIM(L_{1},d)$ ensures that the input data dimensions are consistent with the input layer, $ActivationFeasible(\Phi o,\Phi t)$ ensures that for every $a_{i}\in A$, $Range(X)\subseteq Domain(a_{i})$, i.e., the input data lies within the domain of the activation function, and $GradientExists(\Phi t)$ ensures the existence of a gradient $\nabla_{\theta}L(\theta,X,y)$, i.e., that the loss function $L$ is differentiable with respect to $\theta$.
The learning function $f_{learn}^{2}$ is defined as:
$f_{learn}^{2}(\Phi o,\Phi t)=\{Params(\theta_{new})\land Architecture(L_{1},\ldots,L_{n})\land Activation(A)\}$ (3.10)
where $\theta_{new}=\theta-\eta(\nabla_{\theta}L(\theta,X,y)+\beta v_{prev})$ is the iterative update formula for the neural network model.
Consequently, in the updated information $Iu=(ou,Tu_{h},S_{ou},Mn,Tu_{m},S_{cu})$ of the neural network model, the ontology $ou$ is the union of the designer’s conceptualization and the ontology $ot$ of the training information, $Tu_{h}$ is the set of training times for $Mn$,
$S_{ou}(ou,Tu_{h})=f_{learn}^{2}(\Phi o,\Phi t)$ (3.11)
constitutes the optimized ontological state after learning, with the carrier of $Iu$ being $Mn$ itself, the reflection time $Tu_{m}$ being the post-training period, and $S_{cu}(Mn,Tu_{m})$ as the physical realization of $S_{ou}(ou,Tu_{h})$ within $Mn$, ideally satisfying the mathematical and physical requirements of the enabling mapping:
$S_{cu}(Mn,Tu_{m})=f_{learn}^{2}(\Psi o,\Psi t)$ (3.12)
where $\Psi o$ and $\Psi t$ are the sets of well-formed formulas corresponding to the carrier states of $Io$ and $It$, that is, the physical realizations of $\Phi o$ and $\Phi t$ within $Mn$ and its interface. Equations (3.11) and (3.12) together demonstrate that, within $Mn$, both the ontological state $S_{oo}(oo,To_{h})$ and the carrier state $S_{co}(co,To_{m})$ possess the capability to learn from $S_{ot}(ot,Tt_{h})$ and $S_{ct}(ct,Tt_{m})$, and then optimize to $S_{ou}(ou,Tu_{h})$ and $S_{cu}(cu,Tu_{m})$, respectively; thus, $Mn$ indeed realizes the learning and optimization of $Io$ with respect to $It$, resulting in $Iu$. This completes the proof.
It should be noted that some of the content in the above discussion does not strictly adopt set definitions and well-formed formulas. This does not mean that they cannot be converted into complete set definitions and well-formed formulas, but rather because such representation is more concise and intuitive, and does not lead to ambiguity.

### v9

@@ 2.6 Noisy Information
In fact, for both information $I$ and its noisy information $\tilde{I}$, it is easy to verify its recoverability or recoverable reducibility from the perspective of enabling mapping. However, in real scenarios, it is often difficult to obtain noise-free information $I$, and one can only study using a noisy version $\tilde{I}$. Moreover, the recoverable state of $\tilde{I}$ is not the same as that of $I$, which is the fundamental reason for the skepticism about the recoverability of information. But under the condition where the noiseless information IIcannot be directly obtained, we can approximate the ideal state of $I$ using $\tilde{I}$ along with the relationships between $S_{o}(o,T_{h}),S^{-}_{o}(o,T_{h})$, and $S_{o}^{+}(o,T_{h})$. This is the practical significance of introducing the concept of noisy information.
Shannon also started by discussing the ideal noiseless channel and then discussed the more complicated noisy channel scenarioShannon1948 . In this case, Shannonassumes that the signal generated by the source is $S$, the channel noise is $N$, and the channel output is:
@@
[…]
In this way, we use the definition of noisy information to explain Shannon’s model of noisy channels inShannon1948 , demonstrating that OIT is entirely applicable to the complex scenarios of information research under noisy conditions.
@@ 3 Meta-Framework for Machine Learning Theory (MLT-MF)
The core of machine learning is the artificial intelligence model, which can be regarded as a finite automaton as well as an information system. The processes of operation, learning, testing, and application for such a model all fundamentally depend on information within the system. Therefore, by employing the concepts of finite automaton state and information models established in Section 2, the process of machine learning and its applications can be unified into five primary steps of information realization and functioning:
3.1 Construction of Initial Information
Suppose the model under study is denoted as M, and let $T=\{t_{i}|i=1,\ldots,n\}$ be the set of time points at which M, as a finite automaton, undergoes assignment, input, output, and state transitions in increasing order. According to the perspective of OIT, the creation of M is essentially the realization of information
$I_{o}=\langle M_{D},T_{o_{h}},S_{M_{D}},M,t_{1},S_{M}\rangle$ (3.1)
where the ontology $M_{D}$ of $I_{o}$ is the designer’s conceptualization of M; $T_{o_{h}}$ is the time at which the designer conceives M, and $S_{M_{D}}(M_{D},T_{o_{h}})$ is the initial finite automaton state that must be realized at this conceptual stage. The carrier M of $I_{o}$ is the model itself; the reflection time $t_{1}$ is when M is constructed according to $S_{M_{D}}(M_{D},T_{o_{h}})$; and $S_{M}(M,t_{1})$ is the formalization of the actual physical state of M upon its completion. Thus, $\sup(T_{o_{h}})<\sup(t_{1})$.
According to Theorem 2, let the binary function $\Phi^{2}(X,t)$ represent the set of structures, parameters, and algorithms of model $X$ at time $t$. The formula
$\varphi_{o_{1}}=State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))$ (3.2)
denotes that, at time $t_{1}$, the initial structure, parameters, and algorithm set $\Phi^{2}(M_{D},T_{o_{h}})$ as envisioned by the designer are assigned to $M_{D}$.
$\varphi_{o_{2}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))\land Input^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
$Learnable^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$
$(Eq^{3}(\Phi^{2}(M_{D},t_{2}),f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))))\land State^{3}(M_{D},t_{2},\Phi^{2}(M_{D},t_{2}))$ (3.3)
expresses that, at time $t_{1}$, when the structure, parameters, and algorithm set of $M_{D}$ is $\Phi^{2}(M_{D},T_{o_{h}})$, and the input carrier state $S_{M_{I}}(M_{I},t_{1})$ is provided and satisfies the learnability condition determined by the binary predicate $Learnable^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$, $M_{D}$ can optimize its internal structures, parameters, and algorithmsto $\Phi^{2}(M_{D},t_{2})$ through the learning binary function $f_{learn}^{2}(M_{D},S_{M_{I}}(M_{I},t_{1}))$, assigning this to $M_{D}$. This demonstrates the model’s capacity for learning and optimization, where $M_{I}$ denotes the input device of M.
$\varphi_{o_{3}}=$ $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))\land Input^{3}(M_{D},t_{1},S_{M_{I}}(M_{I},t_{1}))\land$
$Processable^{2}(M,S_{M_{I}}(M_{I},t_{1}))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1}))$
$(Eq^{3}(S_{M_{O}}(M_{O},t_{2}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1})))\land Output^{3}(M,t_{2},S_{M_{O}}(M_{O},t_{2})))$ (3.4)
indicates that, at time $t_{1}$, if $S_{M_{I}}(M_{I},t_{1})$ is input into $M_{D}$ and meets the processability condition specified by the binary predicate $Processable^{2}(M,S_{M_{I}}(M_{I},t_{1}))$, then $M_{D}$ can process $S_{M_{I}}(M_{I},t_{1})$ via the binary processing function $f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{1}))$ to yield the output carrier state $S_{M_{O}}(M_{O},t_{2})$, which is then output at time $t_{2}$. This demonstrates the model’s ability to process and output information, where $M_{O}$ denotes the output device of M.
Thus,
$S_{M_{D}}(M_{D},T_{o_{h}})=\{\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}\}$ (3.5)
constitutes the initial state of M as conceived by the designer.
It should be noted that, since $M_{D}$, $\Phi^{2}(M_{D},T_{o_{h}})$, and $State^{3}(M_{D},t_{1},\Phi^{2}(M_{D},T_{o_{h}}))$ all represent sets, $\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}$ may be considered as three well-formed formulas, or—when decomposed with respect to the elements of each set—as three sets of more granular well-formed formulas. Regardless of the situation, both cases satisfy Definition 4 concerning finite automaton states. Therefore, subsequent discussions will encompass both cases without further elaboration.
$S_{M_{D}}(M_{D},T_{o_{h}})$ may represent the designer’s idea, a design draft on paper, or an electronic document. Through the information $I_{O}$, these are enabled and mapped into the physically operational state of the model M:
$S_{M}(M,t_{1})=\{\psi_{o_{1}},\psi_{o_{2}},\psi_{o_{3}}\}$ (3.6)
where $\psi_{o_{1}}\underline{\triangleleft}I_{o}(\varphi_{o_{1}})$, $\psi_{o_{2}}\underline{\triangleleft}I_{o}(\varphi_{o_{2}})$, and $\psi_{o_{3}}\underline{\triangleleft}I_{o}(\varphi_{o_{3}})$ are the enabling mappings of $\varphi_{o_{1}},\varphi_{o_{2}},\varphi_{o_{3}}$, respectively. This process establishes M as a typical finite automaton and an operational information system.
@@ 3.2 Formal Modeling and Realization of Training Information
According to the general formal definition of information, the formal expression for the realization of training information is straightforward:
$I_{t}=\langle ot,T_{th},S_{ot},M_{I},t_{1},S_{M_{I}}\rangle$ (3.7)
$S_{ot}(ot,T_{th})=\{\varphi_{t}|\varphi_{t}\text{ is a WFF about }ot\text{ and }T_{th}\}$ (3.8)
$S_{M_{I}}(M_{I},t_{1})=\{\psi_{t}|\psi_{t}\text{ is a WFF about }M_{I}\text{ and }t_{1}\}$ (3.9)
Here, the ontology $ot$ refers to the object containing knowledge or samples; $T_{th}$ is the time at which its state is formed; and $S_{ot}(ot,T_{th})$ is a set of business-oriented semantic assertions or facts about the ontology $M_{I}$ functions similarly to the hardware interface and API software of M, carrying the training information at time $t_{1}$. The carrier state $S_{M_{I}}(M_{I},t_{1})$ is the physical realization of $S_{ot}(ot,T_{th})$ via an enabling mapping, ensuring that the information can be machine-read and entered into M.
@@ 3.3 Formal Modeling and Implementation of Model Learning
Model learning for M involves fusing its original information $I_{o}$ with the training information $I_{t}$, resulting in updated information:
$I_{u}=\langle M\cup M_{I},t_{1}\cup t_{2},S_{M\cup M_{I}},M,t_{2},S_{M}\rangle$ (3.10)
Here, the ontology is the hybrid $M\cup M_{I}$, encompassing both the model and its interface; $t_{1}\cup t_{2}$ marks the time of the first learning event; and $S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})$ formally represents the hybrid state during learning:
$S_{M\cup M_{I}}(M\cup M_{I},t_{1}\cup t_{2})=\{\varphi_{u_{1}},\varphi_{u_{2}},\varphi_{u_{3}}\}$ (3.11)
$\varphi_{u_{1}}=Eq^{3}(\Phi^{2}(M,t_{2}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{1})))\land State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))$ (3.12)
indicates that the input state $S_{M_{I}}(M_{I},t_{1})$, which was learned by M at time $t_{1}$, has been transformed into a new set of structures, parameters, and algorithms $\Phi^{2}(M,t_{2})$ at time $t_{2}$, thus becoming the state of M at $t_{2}$.
$\varphi_{u_{2}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
$Learnable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
$(Eq^{3}(\Phi^{2}(M,t_{3}),f_{learn}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land State^{3}(M,t_{3},\Phi^{2}(M,t_{3})))$ (3.13)
shows that after the first learning step, M remains capable of further learning and optimization.
$\varphi_{u_{3}}=$ $State^{3}(M,t_{2},\Phi^{2}(M,t_{2}))\land Input^{3}(M,t_{2},S_{M_{I}}(M_{I},t_{2}))\land$
$Processable^{2}(M,S_{M_{I}}(M_{I},t_{2}))\rightarrow\exists f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2}))$
$(Eq^{3}(S_{M_{O}}(M_{O},t_{3}),f_{process}^{2}(M,S_{M_{I}}(M_{I},t_{2})))\land Output^{3}(M,t_{3},S_{M_{O}}(M_{O},t_{3})))$ (3.14)
indicates that when the training information $I_{t}$ is processable by M, it can also produce output information.
This state is then enabled and mapped to the physical state of the carrier M at time $t_{2}$:
$S_{M}(M,t_{2})=\{\psi_{u_{1}},\psi_{u_{2}},\psi_{u_{3}}\}$ (3.15)